OpenStudy (anonymous):

HELP!!! when a person coughs, the trachea(windpipe contracts, allowing air to be expelled at max velocity. it can be shown that during a cough the velocity v of airflow is given by the function v=f(r)=kx^2(R-r) where r is the radius of the trachea in cm during a cough, R is the trachea's normal radius in cm, and k is a positive constant that depends on the length of the trachea. find the radius r for which the velocity of airflow is greatest.

OpenStudy (anonymous):

Hey ask it in other greoups

OpenStudy (anonymous):

which group??

OpenStudy (anonymous):

this is a maths problem.

OpenStudy (anonymous):

science, biology,

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i think u have to differentiate f(r) to have df(r)/dr, but i can't see what's ment by the "x" in f(r)?

OpenStudy (anonymous):

sorry...thats "r". v=f(r)=kr^2(R-r)

OpenStudy (anonymous):

aha, then: v=kr^2R-kr^2r, where k, R are constants dv/dr=2kRr-3kr^2 finding the maxima by equaling to zero 0=2kRr-3kr^2

OpenStudy (anonymous):

any particular reason for finding the maxima??